OpenStudy (owlcoffee):

Small math question:

OpenStudy (owlcoffee):

Let x,y \(\in \mathbb{R} / x<y\) prove that the following affirmations are true: (1) In the density of \(\mathbb{Q} \) in \(\mathbb{R} \) there exists \(a \in \mathbb{Q} \) such that \(x<a<y\) (2) In the density \(\mathbb{R} / \mathbb{Q} \) in \(\mathbb{R} \) there exists \(\alpha \in \mathbb{R} / \mathbb{Q} \) such that \(x< \alpha < y\)

OpenStudy (anonymous):

Im thinking if we should let \[a=\frac{ x+y }{ 2 }\]

OpenStudy (anonymous):

Hmm Let \[a=\frac{ x+y }{ 2 } {\space}, where {\space} a \in Q\] Now, I'm sure if you want to prove this, then we can assume other simpler theorems to apply. So we can use the fact that \[x+y\] is rational and that the product of rationals \[\left( x+y, {\space} \frac{ 1 }{ 2 } \right)\] is also rational. Also, since it states that, \[x<y\] then adding x to both sides means \[2x<x+y\] And similarly, by adding y to both sides we get: \[x+y<2y\] So therefore we can bound these restrictions essentially into one domain: \[2x<x+y<2y\] This implies that \[x<\frac{ x+y }{ 2 }<y\] , where \[a=\frac{ x+y }{ 2 }\]

OpenStudy (anonymous):

now for part 2, its probs a bit tricky

OpenStudy (anonymous):

perhaps set \[a=\frac{ x+y }{ \sqrt2 }\] Now one tricky bit here is that we can make this value rational if \[x=-y\] which means with some algebra, we get \[a=0\] So consider two cases: Either, \[x \ge 0\] or \[y \le0\] so we can take \[a=\frac{ x+y }{ \sqrt2 }\] whereas if \[x<0<y\], then we can have it set to \[a=\frac{ y }{ \sqrt2 }\] and by now we can see 'a' is necessarily irrational and therefore proof is complete since we can assume an earlier theorem where \[If {\space} x \in Q {\space} and {\space} y \in R/Q, {\space} then {\space} x+y \in R/Q\]

OpenStudy (anonymous):

Probably a good idea to use that assumption above as well as \[If {\space} x \in Q {\space} and {\space} y \in R/Q, {\space} then {\space} xy \in R/Q {\space} provided {\space} x \neq0\]

OpenStudy (anonymous):

sorry if this is a bit general, my real analysis knowledge is a bit distant now

OpenStudy (anonymous):

probs best if i stick with studying my chem eng courses atm haha

OpenStudy (owlcoffee):

Well, this looks good but thing is bounding out the domain would cause troubles when wanting to actually analyze the density of Q in R... I was thinking about calling "n" : \(n \in \mathbb{N} \) such that \(\frac{ 1 }{ n }< y-x \), this will generate a rational, bounded to Q in R such that calling it "a" would be: \[a:=\frac{ E(nx)+1 }{ n }\] And this should work smoothly since "E" would be the whole function defined by: \[E(x):=\max \left\{ p \in \mathbb{Z} :p \le x \right\}\] Now this should be enough to constate that inside these conditions \(x<a<y \) should be verified because of E(x) and it's behaviour. I have to research that on the library in the afternoon, but thanks a lot for giving this question some time haha.

OpenStudy (anonymous):

yes very curious to see!

OpenStudy (anonymous):

let me try and think maybe haha good though process though

OpenStudy (owlcoffee):

I based my conclusion on: \[\frac{ E(nx) }{ n } \le x< \frac{ E(nx)+1 }{ n }\] \[x<a \le x+\frac{ 1 }{ n } < x+(y-x)=y\]

OpenStudy (anonymous):

interesting

OpenStudy (anonymous):

im kinda stuck now haha

OpenStudy (owlcoffee):

Haha, I am also a lil' stuck but I think it looks right, I am also quite rusty when it comes to real analysis.

OpenStudy (anonymous):

perhaps share me the verdict when you find out? I wouldn't mind delving back into real analysis for a bit haha

OpenStudy (anonymous):

good question though! i'm off to bed. catchya